Dalton's age is half of Chen's. In 10 years Chen's age will be three times what Dalton's was 5 years ago. What are their ages?
Answer by mananth(16946)

Dalton's age is half of Chen's. In 10 years Chen's age will be three times what Dalton's was 5 years ago. What are their ages?
...
let Chen's age be x
Dalton's age = x/2
..
after 10 years
chen's age = x+10
Dalton's age = (x/2) +10
..
five years ago Dalton was (x/2)-5 years
..
x+10= 3((x/2)-5)
x+10=3(x-10)/2
x+10=3x-30/2
multiply by 2
2(x+10)=3x-30
2x+20=3x-30
2x-3x=-50
-x=-50
x= 50 chen's age
Dalton's age = x/2 = 25 years
...
m.ananth@hotmail.ca